Hi,
Can you please help me plot the graph of :
It is suppose to be a linear equation in one variable
2x+3=4

I plotted it but I am not sure since I don't know how to graph.
Thank you,
Mark

Answer by checkley77(12569)

2X+3=4
2X=4-3
2X=1
X=1/2=.5 THIS IS A VERTICLE LINE THROUGH THE POINT (0.5,0)